⟨SK_{i}, P K_{i}⟩, P K_{r} i

⟨SK_{r}, P K_{r}⟩, P K_{i}

r

nonce ni

{i, ni}P K_{r}
nonce nj
{ni, nr}P K_{i}
{nr}P K_{r}

i and r are communicating i and r share secrets ni, nj msc Needham-Schroeder Protocol